A sample of 20 account balances of a credit company showed an average balance of $1,180 and a standard deviation of $125. You want to determine if the mean of all account balances is significantly greater than $1,150. Assume the population of account balances is normally distributed. Compute the p-value for this test.
Answer #1

Here we have sample size (n) = 20 Sample mean W = 1180 Standard deviationo)=125 Stephs Hypothesis Ho 4: 1150 Hai l> 1150 - T

Step 3: Pralue Pralce at z = 11.071 Rvalue = P1Z) 1.07) = 1- P1Z5 1.07 = 1-0.857691 from e tandard | normal calculator = 0.14

Dear student, please like it.

Thanks.